two string are joined .the linear mass density of string 2 is 4 times string 1.the incident wave is partially reflected and partially transmitted.incident wave equation is yi=Ai sin(k1x-w1t),find reflected and trasmitted wave equation

Grade:12

1 Answers

Sher Mohammad IIT Delhi
9 years ago
the coefficient of transmission = 4*sqrt(1*4)/(1+42)= 8/17
the coefficient of reflection =1- 8/17= 9/17
equation of transmitted wave = 8/17 Ai sin(k1x-w1t)
equation of reflected wave = 9/17 Ai sin(k1x+w1t)
sher mohammad
askiitians faculty
b.tech, iit delhi